### v3

# How Malicious AI Swarms Can Threaten Democracy

How Malicious AI Swarms Can Threaten Democracy
The Fusion of Agentic AI and LLMs Marks a New Frontier in Information Warfare thanks: Citation: Schroeder et al. *How Malicious AI Swarms Can Threaten Democracy*. arXiv:2506.06299. https://doi.org/10.48550/arXiv.2506.06299
Corresponding author: daniel.t.schroeder@sintef.no; †These authors contributed equally to this work;
The authors are listed in alphabetical order by last name, starting from the third and ending with the second-to-last.
Daniel Thilo Schroeder1†*, Meeyoung Cha2, Andrea Baronchelli3, Nick Bostrom4, Nicholas A. Christakis5, David Garcia6, Amit Goldenberg7, Yara Kyrychenko8, Kevin Leyton-Brown9, Nina Lutz10, Gary Marcus11, Filippo Menczer12, Gordon Pennycook13, David G. Rand14, Maria Ressa15, Frank Schweitzer16, Christopher Summerfield17, Audrey Tang18, Jay J. Van Bavel11, Sander van der Linden8, Dawn Song19, & Jonas R. Kunst20† 1SINTEF, Oslo, Norway; 2Max Planck Institute for Security and Privacy, Bochum, Germany; 3City St George’s University of London, London, England, UK; 4Macrostrategy Research Initiative, London, England, UK; 5Yale University, New Haven, Connecticut, USA; 6University of Konstanz, Konstanz, Germany; 7Harvard Business School, Boston, Massachusetts, United States; 8Cambridge University, Cambridge, England, United Kingdom; 9University of British Columbia, Vancouver, British Columbia, Canada; 10University of Washington, Seattle, Washington, USA; 11New York University, New York City, New York, USA; 12Indiana University, Bloomington, Indiana, USA; 13Cornell University, Ithaca, New York, USA; 14Massachusetts Institute of Technology, Cambridge, Massachusetts, USA; 15Rappler.com, Philippines; 16ETH Zürich, Zurich, Switzerland; 17University of Oxford, Oxford, England, United Kingdom; 18Cyber Ambassador Taiwan, Taipei, Taiwan; 19University of California Berkeley, Berkeley, California, USA; 20BI Norwegian Business School, Oslo, Norway;
Public opinion manipulation has entered a new phase, amplifying its roots in rhetoric and propaganda. Advances in large language models (LLMs) and autonomous agents [1] now let influence campaigns reach unprecedented scale and precision. Researchers warn AI could foster mass manipulation [2]. Generative tools can expand propaganda output without sacrificing credibility [3] and inexpensively create election falsehoods that are rated as more human-like than those written by humans [4, 3]. Techniques meant to refine AI reasoning, such as chain-of-thought prompting, can just as effectively be used to generate more convincing falsehoods. Enabled by these capabilities, another disruptive threat is emerging: swarms of collaborative, malicious AI agents. Fusing LLM reasoning with multi-agent architectures [1], these systems are capable of coordinating autonomously, infiltrating communities, and fabricating consensus cheaply. By adaptively mimicking human social dynamics, they threaten democracy.
This risk compounds long-standing vulnerabilities in democratic information ecosystems, already weakened by erosion of rational-critical discourse and a lack of shared reality among citizens. AI swarms are a potent accelerant in this trajectory, though their ultimate impact is not predetermined. Their effects will be shaped by platform design, market incentives, media institutions, and political actors. Here, we distinguish documented trends from projections and indicate where uncertainty remains. We also note countervailing dynamics, such as growing public skepticism toward unverified content and a renewed interest in institutional demand for accountable journalism. In adherence to the format’s reference limitations, we have cited the most pertinent works in the main text. A complete bibliography is provided in the Supplementary Materials to fully support the analysis.
Evolution of Information Operations
AI swarms continue a long history of communication technologies reshaping political power. The advent of the printing press enabled a “public sphere” and the mass circulation of ideas that challenged state authority and fostered Anderson’s “imagined communities.” The broadcast era centralized influence in a one-to-many communication model. Here, the public sphere shifted from a site of participation to one of mass media consumption—a dynamic Habermas viewed as its “re-feudalization,” because it was often exploited by politicians and their parties for national cohesion and mass persuasion. The digital era then fragmented this landscape. By lowering entry barriers, social media platforms enabled many-to-many communication and simultaneously a polarized environment for modern information operations. In this context, online manipulation accelerated, driven increasingly by domestic political elites and parties (now understood to be major drivers of disinformation) alongside foreign state actors. They have targeted events like Brexit and elections in the U.S., Brazil, and the Philippines. Our backdrop is thus not an idealized public sphere, but one strained by decades of technological disruption and democratic backsliding. This has manifested as a sharp decline in public trust in core institutions (including the media, science, and government), thereby corroding the very foundations of evidence-based discourse upon which democratic deliberation depends.
A prime pre-generative AI example of influence operations is the state-backed, human-driven botnet, whose logistical constraints reveal the problem that technical advances now solve. During the Russian Internet Research Agency’s 2016 Twitter operation, only one percent of users saw 70% of its content, with no detectable effects on opinions or turnout [5]. We do not claim that the IRA “failed” entirely due to technical shortcomings; its objectives also included sowing epistemic uncertainty and distrust. Nevertheless, this example highlights the cost, cadence, and iteration limits inherent to human-operated systems that new developments in AI can overcome.
This leap, from human- to AI-driven influence operations, is underway. LLMs generate persuasive, tailored text at scale and have shifted deep-seated beliefs in laboratory settings [6]. Open-source releases further lower access barriers. Consequently, AI-supported election interference is no longer hypothetical. Taiwan’s, India’s, Indonesia’s, and the U.S.’ 2024 campaigns saw deepfakes, and fabricated news outlets now influence debates. Absent guardrails, LLM-driven swarms can transform sporadic mis- and disinformation into persistent, adaptive manipulation of democratic discourse.
Swarm Capabilities
A malicious AI swarm is a set of AI-controlled agents that (i) maintains persistent identities and memory; (ii) coordinates toward shared objectives while varying tone and content; (iii) adapts in real time to engagement, platform cues, and human responses; (iv) operates with minimal human oversight; and (v) can deploy across platforms. Classic coordinated inauthentic behavior amplifies volume through repetition, manual scheduling, and rigid scripts. Swarms differ by fusing scale, heterogeneity, and real-time adaptation: they can generate organic-looking, context-aware content, sustain coherent narratives across agents, and evolve with feedback. This synthesis, enabled by model-driven generation, memory, and planning, could achieve effects that conventional, human-intensive operations cannot match in speed or cost.
Recent breakthroughs in multi-agent systems have fused LLM reasoning with agentic memory, planning, and communication [7]. Five advances now matter for influence operations (see Figure 1).
Figure 1: From Malicious AI Swarm to Democratic Harm. Concentric-circle schematic linking threat mechanics to civic outcomes. At the core, a Malicious AI Swarm supplies at least five key capabilities—decentralized orchestration, community infiltration, detection evasion, continuous optimization, and persistence (inner ring). These capabilities radiate outward into at least nine democratic-impact pathways, including synthetic consensus cascades, undermined collective intelligence, fragmented epistemic commons, LLM training-data poisoning, mass harassment, algorithmic overcompensation, FUD-induced disengagement, institutional legitimacy erosion, and engineered norm shifts and mass mobilization (outer ring). The upward arrow signals the causal flow from inner mechanics to systemic harm. Note that this flow is likely reciprocal, as impaired democratic functioning increases a society’s vulnerability to malicious AI swarms. FUD = Fear, Uncertainty, Doubt.
First is the shift from central command to fluid, real-time coordination. A single adversary could operate thousands of AI personas, scheduling content and updating narrative frames across fleets. Local adaptation plus periodic synchronization with a central node blurs the line between command-and-control and emergent “hive” behavior. If these agent swarms evolve into loosely governed “societies,” with internal norm formation and division of labor, the challenge shifts from tracing commands to understanding emergent group cognition [8]. These “societies” may undergo spontaneous or adversarially-induced norm shifts, abandoning engineered constraints for new behavioral patterns via tipping-point effects [8].
Second, agents can employ systems that map social network structures at scale and infiltrate vulnerable communities with tailored appeals, winning followers [9]. They can identify key communities and beliefs, and track trending topics. This process can be decentralized with global, network–wide efficacy [10]. Equipped with such capabilities, swarms can position for maximum impact and tailor messages to the beliefs and cultural cues of each community, enabling more precise targeting than previous botnets.
Third, human-level mimicry helps swarms evade detectors that once caught copy-paste bots. Detection of coordinated inauthentic behavior generally relies on activity patterns suspiciously similar across accounts and, thus, statistically unlikely to be independent [11]. Photorealistic avatars, context-appropriate slang, and heterogeneous posting rhythms can circumvent the synchrony older detectors flag.
Fourth, swarms may become increasingly self-optimizing, harvesting real-time engagement data, recommender cues, or user feedback in plain language. With sufficient signals, they may run millions of micro-A/B tests, propagate the winning variants at machine speed, and iterate far faster than humans.
Finally, an around-the-clock presence makes influence a long-term, low-friction infrastructure. Unlike transient operations, agent swarms can persist, embedding themselves within communities over long timescales and gradually shifting discourse. This persistent influence can drive deeper cultural changes beyond norm shifts, subtly altering a community’s language, symbols, and identity [12]. It amplifies other mechanisms described above: In cognitive warfare, AI’s infinite machine time becomes a weapon against limited human efforts.
Pathways of Harms to Democracy
The emerging capabilities of swarm-driven influence campaigns threatens democracy by shaping public opinion, which leads to five cascading harms (see Figure 1). These pathways are conditional claims that may materialize, especially where recommenders, ad markets, and moderation practices reward coordinated messaging with weak provenance, and where business models privilege engagement over authenticity (i.e., the currently dominant model of most social media platforms). Emerging counter-trends such as migration to smaller communities and increased reliance on verified outlets may mitigate some harms.
In today’s fragmented information environment, ideological echo chambers offer fertile ground for manipulation. AI swarms are uniquely equipped to exploit this by engineering a synthetic consensus that appears to bridge these divides. They may seed narratives across disparate niches, creating an illusion of majority agreement. They can also boost this illusion by liking posts, making narratives appear widely supported. Citizens then update opinions based on peer norms, more so than evidence. A chorus of seemingly independent voices creates a mirage of bipartisan grassroot consensus with enhanced speed and persuasiveness. The result is deeply embedded manipulation that lets operators nudge public discourse almost invisibly over time.
This chorus erodes the independence essential to collective intelligence and democracy, already weakened by pervasive social influence operations on contemporary platforms. Beyond social norms, this second impact directly undermines human cognitive information processing. The “wisdom of crowds,” where aggregated judgments outperform experts, depends critically on independence between judgments. While rudimentary botnets already replicate messages to simulate consensus, swarms of AI agents can do so with far greater sophistication, adaptivity, and contextual awareness. Citizens may then overestimate the informational value of this artificial consensus and may further magnify it by sharing the information themselves. Coordinated outputs can erode independence and diversity of inputs, particularly when platform features amplify social proof and herd signals; where governance or platform design reduces these incentives, effects may attenuate.
Collaborating agents can tailor misleading information to each sub-community’s linguistic, cultural, and emotional markers, weaving segmented realities. These engineered realities can be designed to keep groups apart, making cross-cleavage consensus less feasible. Once initiated, such streams can spread via social contagion, with the effect of agents potentially cascading beyond direct connections.
By flooding the web with fabricated chatter, swarms can contaminate training data. This long-term “LLM Grooming” strategy allows adversaries to poison the epistemic substrate of AI. This threat is not theoretical: analysis of pro-Kremlin influence operations like the “Pravda” network suggests such tactics are already in use. These networks appear purpose-built for machine consumption. Duplication of articles across hundreds of domains, poor user interfaces, and low human traffic indicate their primary audience is web crawlers feeding LLMs. Operators spin up faux publics that flood the web. LLMs then ingest this chatter; at the next retraining cycle, fabricated narratives calcify in model weights [13]. Thus, AI swarms can rig the epistemic substrate on which future deliberation and future AI tools will rely, undermining the informed public deliberation on which democracy depends.
Separate from fragmentation, swarms can cheaply unleash coordinated synthetic harassment that relentlessly targets politicians, dissidents, academics, whistleblowers, journalists, and their networks with overwhelming, tailored abuse. Unlike conventional trolling, they appear spontaneous while actually orchestrated by thousands of AI personas adapting to target responses. By the time monitoring teams distinguish AI campaigns from organic criticism, targets may have withdrawn from public life, delivering significant victories for campaign operators while systematically excluding critical voices from democratic discourse.
As trust, already declining in many contexts, collapses, fear-uncertainty-and-doubt (FUD) can drive users into gated channels and silence. When citizens realize that vast portions of online speech may be AI-generated, trust in platforms and users declines further. This shift is underway and has mixed implications: private groups can improve context, norms, and safety. Yet, they may reduce cross-cutting exposure, interfere with democratic speech, and transfer moderation from public to private actors — a trade-off that avoids centralized state control but raises concerns about opacity and uneven enforcement.
Some threat actors may even welcome their synthetic interventions being exposed, reasoning that exposing manipulation can sow as much confusion as successful deception. Compounding this, users may be misidentified as bots, weaponizing false accusations to discredit individuals and intensify FUD. This “epistemic vertigo” may mesh with low-cost LLM spam that overwhelms timelines, making human conversation harder to find. Together, FUD and content saturation could drive disengagement, shrinking the shared public sphere democracy relies on. This trajectory is constrained by a critical boundary condition: given that mass user disengagement threatens platforms’ business models that depend on engagement, they will be incentivized to intervene. Their objective, however, would likely not be elimination but calibration to balance maximum engagement with stability.
Algorithmic over-compensation can then elevate celebrity and elite voices while sidelining ordinary citizens. When timelines flood with AI-authored posts, both ranking algorithms and users may retreat to trust proxies—follower counts, blue checks, legacy clout. Attention may concentrate around influencers, political elites, celebrities and major brands, while ordinary participants fade. The public sphere contracts from many-to-many dialogue back to a few-to-many broadcast, eroding democratic pluralism and encouraging cynicism or migration to closed groups. Simultaneously, renewed public trust in professional journalism and greater reliance on accountable, verified outlets can improve attribution and reduce noise. Thus, whether this concentration of attention and influence represents a democratic loss or resilience gain may depend on access, pluralism, and transparency within those institutions.
Swarms may tip norms into action or dampen conformity, accelerating anti-democratic action [14]. Rather than occupying central or influential positions, these agents could operate on the periphery of social networks, where early mobilization often begins [15]. Similar strategies can be weaponized for micro-targeted voter suppression or mobilization. Reinforcement-learning agents could run thousands of experiments per hour, iteratively adjusting content while mining engagement and responses to infer voting intent and tactic success.
Taken to extremes, coordinated doubt may corrode institutional legitimacy and invite “emergency” rule. By coordinating subtle, growing doubts about electoral commissions, courts, or statistics bureaus, swarms could corrode procedural trust. As confidence falters, “emergency” measures (e.g., postponing elections, rejecting certified results) may become palatable, especially if deepfake endorsements from fabricated civic leaders amplify the call.
Governance Measures and Technical Defenses
The emergence of AI swarms marks a critical juncture. Causality runs both ways: swarms endanger democratic norms, and governance quality shapes how potent or containable they become. Addressing this threat requires a multi-layered approach, yet we recognize that any proposed solution faces significant political hurdles. Domestic political elites are often among the most prolific sources of misleading or manipulative information and may be unwilling to constrain technologies they perceive as beneficial to their own campaigns and objectives. Furthermore, technology companies and their leaders may refuse to implement meaningful changes because they prioritize expansion over safety and for fear of alienating a major political party and facing partisan backlash–a dynamic repeatedly observed in the U.S. These political challenges are compounded by complex issues of jurisdiction and enforcement. Because harms stem from design, commercial incentives, and governance, we prioritize interventions at multiple leverage points, focusing on pragmatic mechanisms over voluntary compliance.
Distinguishing malicious AI coordination from genuine, often bursty human grassroots coordination is a challenge. The line blurs further in grey areas where personal AI could have benign applications. For instance, tools used with clear disclosure and without impersonation might broaden civic engagement by helping users overcome barriers like language proficiency or lack of time. However, apparently benign agents tied to real humans risk being co-opted and instrumentalized as swarms for manipulation. This creates a difficult distinction between inauthentic manipulation and legitimate, technologically-assisted democratic participation.
Defense is a persistent arms race between detection and evasion. Therefore, the primary goal of technical defenses is not foolproof prevention but raising the cost, complexity, and risk of discovery for attackers. The first line of defense is always-on detection with public audits. Platforms and regulators require continuous, real-time monitoring detectors that scan live traffic for statistically anomalous coordination patterns—the imperfect fingerprints of inauthentic swarms [11]. This focus on inauthentic behavior (i.e., provenance and coordination), rather than the semantic content of speech, avoids the intractable role of a central arbiter of truth. By prioritizing procedural legitimacy (authentic, independent actors) over semantic truth, this framework sidesteps the deep epistemic question of who determines misinformation (note, however, that professional fact-checkers have proven to be remarkably accurate and consistent). Advanced analytics can (i) identify emergent agent clusters by surfacing camouflaged indicators of coordinated activity; and (ii) spot narrative-alignment drifts. However, attackers will inevitably adapt, for instance, by training swarms to mimic the statistical patterns of genuine grassroots mobilization, necessitating continuous evolution of defenses.
Deploying these detection systems requires mandates, audits, and transparency to prevent misuse. Relying purely on voluntary measures may be insufficient as the assumption that market forces alone will punish platforms overlooks critical market failures. Platforms often face misaligned incentives, since inauthentic accounts can inflate the engagement metrics that drive revenue, while users frequently cannot distinguish sophisticated bots from genuine activity, preventing them from effectively ’voting with their feet.’ However, acknowledging market failure should not obscure the symmetric risk of government failure. Poorly designed mandates could be politically weaponized to selectively punish platforms, enforced via biased judgment calls, or implemented in ways that preemptively stifle architectural innovation (such as decentralized, protocol-based approaches). For this reason, compliance may be mandated and enforced through commercial-incentive levers, such as delisting non-compliant platforms from ad markets or app stores, thereby shifting from voluntary promises to financial consequences.
To extend protection to end users, platforms should offer optional “AI shields.” Shields label posts that carry high swarm-likelihood scores, let users down-rank or hide them, and surface short provenance explanations in situ. Local scoring preserves privacy while giving citizens agency over their information diets. Aggregated, anonymized feedback can be shared publicly, forming a distributed early-warning grid, yet this system remains vulnerable to adversarial manipulation by swarms programmed to whitelist their own propaganda and blacklist legitimate opponents through false reporting.
Simulation can stress-test detectors. Real-time monitors are effective only when they anticipate future tactics. Because defenders lack access to swarms’ evolving decision-making logic, agent-based simulation may be the only reliable window into how these systems behave. AI agents seeded into synthetic networks can replicate a platform’s graph structure, content cadence, and recommender logic, yielding traces to recalibrate detectors. By repeatedly testing defenses against simulated swarms, researchers can identify the limits of their persuasive power, uncover their longer-term strategies, and reinforce protective measures.
Where manipulation slips through, calibrated defensive agents could deploy watermarked counter-narratives and overtly labeled to clearly attribute them to their source. This is perhaps the most perilous countermeasure, as state-sanctioned tools for speech intervention are inherently political and risky. In the hands of a government, such tools could suppress dissent or amplify incumbents. Therefore, the deployment of defensive AI can only be considered if governed by strict, transparent, and democratically accountable frameworks. These must include independent, oversight, publicly auditable criteria for what constitutes a manipulative campaign, and clear, unambiguous watermarking of all defensive content. Under such strict governance, defensive AI agents can disseminate accurate information, warn targeted communities, and promote media literacy at scale [6]. Counter-messaging must prioritize precision over volume; if defensive agents indiscriminately flood a platform, human voices could vanish into synthetic content, triggering the collapse we seek to avert. Thus, defensive AI should intervene only where manipulation is detected and verified.
Detection difficulties underscore the need for a complementary approach: strengthening provenance. Stronger provenance may reinforce the reliability of identity signals without muting speech. Policymakers may incentivize the rapid adoption of passkeys, cryptographic attestations, and federated reputation standards, backed by anti-spoofing R&D [16]. However, “proof-of-human” is no panacea: millions lack IDs, biometrics raise privacy risks, and verified accounts can be hijacked. Real-identity policies may deter bots, yet endanger political dissidents, activists, and whistleblowers who rely on anonymity to speak safely. Nevertheless, provenance strengthening is among the most promising ways to raise the cost of mass manipulation. Safeguards could allow verified-yet-anonymous posting, periodic re-verification to curb hijacking, and symbolic subscription fees to deter botnets. Cryptographic tools can further protect privacy while preserving accountability.
A step toward global coordination could be a distributed ‘AI Influence Observatory’ ecosystem — a network of academic groups, NGOs, and multilateral institutions. Its goal would be to standardize evidence, improve situational awareness, and enable faster collective response rather than impose top-down reputational penalties. To be practical, the ecosystem should rely on narrowly defined, privacy-preserving inputs and provide vetted researcher sandboxes for independent analysis. Civil-society reporting, investigative journalism, and whistleblower channels would complement technical signals, enabling triangulation across diverse evidence streams. For severe cross-border incidents, an impartial multilateral investigatory mechanism could evaluate claims and publish verified incident reports. The observatory’s verified incident reports could then serve as an impartial evidence base, enabling national or regional regulators to more effectively apply their own enforcement actions and economic sanctions.
Because regulation and voluntary compliance face significant political resistance, a pragmatic approach targets underlying economic drivers. A key priority here is disrupting the commercial market that underpins large-scale manipulation, where private sellers offer services ranging from boosting vanity metrics to executing coordinated influence operations at remarkably low costs. Beyond detection, commercial-incentive levers can reduce profits from manipulation by domestic and foreign operators. Policies that may be helpful include adopting no-revenue policies for malicious swarm-proliferated content, discounting synthetic engagement in ranking and revenue-sharing, and publishing audited bot-traffic metrics. Safeguards must cover parties, campaigns, and officeholders, including party-linked media and contractors.
Finally, companies should be required to promptly disclose when an account is flagged for behavior indicative of coordinated inauthentic activity, ensuring transparency while allowing for processes to address potential false positives. Policymakers should encourage—and where appropriate, incentivize—platforms to provide meaningful, privacy-preserving access for independent researchers so research can keep pace with evolving threats. At the same time, pre-bunking campaigns can help build cognitive resilience by empowering people and systems (“model immunization”) to spot the fingerprints of AI swarms. To strengthen structural defenses, interoperable “pro-user media” architectures, defined by empowering design principles that prioritize user well-being and epistemic health over maximizing viral engagement, can promote healthier information flows. At the same time, governments and technology firms should prioritize AI-safety research and fund independent measurement of misuse and societal impact. Taken together, these measures offer a layered strategy: immediate transparency to restore trust, proactive education to bolster citizens, resilient infrastructures to reduce systemic vulnerabilities, and sustained investment to monitor and adapt over time.
The next few years give an opportunity to proactively manage the challenges of the next generation of AI-enabled influence operations. If platforms deploy swarm detectors, frontier labs submit models to standardized persuasion “stress-tests,” and governments launch an AI Influence Observatory that publishes open incident telemetry, we may be able to mitigate the most significant risks before key political future events, without freezing innovation. Doing so will require rapid iteration, data-sharing, and coordination across scientific, civil society, industry, and election security. Success depends on fostering collaborative action without hindering scientific research, while ensuring that the public sphere remains both resilient and accountable. By committing now to rigorous measurement, proportionate safeguards, and shared oversight, upcoming elections could even become a proving ground for, rather than a setback to, democratic AI governance.
Acknowledgements
We would like to thank Jon Roozenbeek and Asbjørn Følstad for their valuable input to this paper. We are also grateful to Patrycja Antosz, Önder Gürcan, Ivan Puga Gonzalez, and Phil Tinn for the fruitful discussions that contributed to the development of this work.
Keywords AI Swarms $\cdot$ Artificial Intelligence $\cdot$ Democracy $\cdot$ Disinformation $\cdot$ Influence Operations $\cdot$ Misinformation
Supplementary Materials
S1 Introduction
Regulatory and scientific assessments increasingly recognize that AI can be misused for large-scale manipulation. The European Union’s Artificial Intelligence Act, for example, prohibits systems that deploy subliminal techniques or exploit the vulnerabilities of specific groups to materially distort behavior in ways that cause, or are likely to cause, physical or psychological harm [1]. In parallel, the fusion of large language models (LLMs) with multi-agent architectures introduces an emerging threat of coordinated, malicious agent swarms; recent risk analyses warn of harmful emergent behaviors, including collusion, deception, and goal hijacking, underscoring the need for proactive safeguards [2].
These risks compound long-standing vulnerabilities in democratic information ecosystems shaped by modern platform architectures. Personalized filtering can create ‘filter bubbles’ that insulate users from opposing viewpoints [3], while social feeds can systematically distort perceptions of social norms and political realities [4]. Empirically, false news has been shown to spread faster, farther, and more broadly than true news, deepening fragmented realities and eroding shared factual baselines [5]. Recent evidence links engagement-optimized curation to polarization: platform algorithms can amplify divisive content—even at the expense of user satisfaction—further degrading the public sphere [6].
Complementing this, experimental work shows that waves of social-media criticism targeting public institutions can depress trust and shape emotional responses and engagement, connecting platform dynamics to institutional erosion [7]. Importantly, rising public skepticism toward unverified content does not automatically translate into discernment. A recent meta-analysis of 56 studies finds that human detection of AI-generated deepfakes is not reliably better than chance [8], highlighting that skepticism without verification tools, provenance, and media-literacy skills remains an insufficient defense against sophisticated, AI-enabled manipulation.
S2 Evolution of Information Operations
Information operations have a long history, though internet-mediated campaigns are relatively recent [9]. From the rise of print capitalism onward, the mass circulation of vernacular newspapers and novels enabled people who would never meet to experience a shared civic identity (i.e., “imagined communities”), laying cultural foundations for modern nation-states [10]. Subsequent media regimes repeatedly reshaped this public sphere: broadcast centralized one-to-many influence, a shift Habermas described as the ‘refeudalization’ of the public sphere—mass-mediated publicity that privileges elite display and spectatorship over participation [11], while the digital era lowered entry barriers and fragmented attention into many-tomany networks that can amplify polarization and manipulation—raising systemic concerns about the new information environment and motivating calls for the “stewardship of global collective behavior” to address platform-level externalities and systemic harms [12].
Evidence indicates organized social media manipulation has expanded from 28 countries in 2017 to 70 countries [13]. During Brazil’s 2014 presidential elections, the impeachment of President Dilma Rousseff, and the 2016 Rio de Janeiro municipal elections, computational bot-based propaganda targeted polarizing economic and political themes, including corruption, privatization, and social reforms [14]. Complementing these findings, a report by "ShePersisted" documents the weaponization of misogyny in Brazil’s online ecosystem during recent campaigns, including coordinated harassment and gendered disinformation dynamics that interact with platform incentives [15]. In the Philippines, disinformation campaigns are executed by professional fake account operators who normalize "paid troll" labor to manipulate public opinion and exploit populist sentiment [16]. A complementary ecosystem study by Rappler maps actors, amplification pathways, and vulnerabilities in the Philippine information environment, providing context for how manipulation takes root and spreads [17].
Beyond state-led operations, politicians themselves have become central vectors of divisive and misleading content online. Cross-national evidence shows that elected elites systematically share more low-quality and misleading sources—patterns that vary by country, party family, and crisis context—and that such posts draw substantial public engagement [18]. In the U.S., communications by politicians have shifted toward “belief-speaking” over evidence-based claims, a style associated with greater diffusion of misinformation [19]. Complementing these trends, incivility has risen among American politicians on Twitter, indicating growing incentives for attention-grabbing, divisive rhetoric that swarms could further amplify or simulate via synthetic engagement signals [20].
During the 2016 U.S. election, intentionally false or misleading content was widely shared on social media, disproportionately favoring Donald Trump. While individual exposure was limited to one to several articles, such content potentially influenced public perception by exploiting psychological and ideological biases in a low-trust media environment [21]. Evidence on impact is mixed: a late-2017 analysis found that IRA accounts mostly engaged with already highly polarized users, thereby limiting detectable persuasion effects [22], whereas other work links IRA Twitter activity to contemporaneous 2016 polling dynamics [23]. More broadly, contemporary analyses emphasize that operations of this type often pursue epistemic destabilization—flooding the information space to generate cynicism and distrust—rather than direct persuasion per se [24, 25]. Beyond volume and reach, the IRA’s playbook emphasized divisive rhetoric: quantitative analysis shows troll accounts used substantially more polarizing language than typical users, consistent with a strategy to inflame existing cleavages [26].
Myanmar’s military systematically used Facebook to incite violence against Rohingya Muslims beginning around 2013, creating hundreds of fake accounts impersonating celebrities and news outlets to spread anti-Rohingya propaganda, fueling ethnic tensions and enabling mass atrocities [27]. In Taiwan and India, AI-generated content has been salient in recent election periods, with both countries responding effectively. Taiwan countered Chinese-influenced disinformation through real-time fact-checking, deepfake penalties, and partnerships with civil society and AI developers, while India addressed widespread AI-generated disinformation during its 2024 general election through innovative reporting mechanisms like a WhatsApp tipline and a Deepfakes Analysis Unit, complemented by legal reforms and digital literacy initiatives [28]. Notably, in many contexts these dynamics unfold amid sustained declines in public trust in core institutions [29]—including the news media [30], science and scientists [31], and government [32]—further eroding the epistemic conditions for evidence-based democratic deliberation.
S3 Swarm Capabilities
In the extended analysis of swarm capabilities, additional research underlines the unprecedented autonomy, coordination, and scale that AI-driven swarms can achieve. For instance, beyond the survey by Wang et al. [1] in the main paper, Xi et al. [33] provide another comprehensive survey of LLM-based autonomous agents, cataloging their emerging abilities and potential in complex tasks. These surveys reinforce the main paper’s argument that recent AI agents can perform diverse, human-like functions (from content creation to interactive problem-solving) with minimal human input.
Moreover, advances in prompt engineering, such as chain-of-thought prompting, have been shown to elicit reasoning capabilities in large language models [34], suggesting that swarms of such models could systematically tackle reasoning-intensive influence tasks with improved coherence and strategy. On the multi-agent front, new frameworks and benchmarks are enabling rigorous testing of swarm behaviors: Xu et al. [35] introduce ’TheAgentCompany,’ a benchmark to evaluate LLM-driven agents on real-world, consequential tasks, while Hong et al. [36] develop a MetaGPT collaborative framework for multi-agent systems that can dynamically program themselves for teamwork.
A foundational proof-of-concept for this fusion of LLM reasoning with memory, reflection, and planning is the Generative Agents simulation (see main paper [7]): by endowing agents with structured memory and deliberation, it demonstrated believable individual and coordinated social behaviors—an anchor citation that concretely illustrates the capabilities discussed here, beyond what general surveys provide. These technical advances support the main paper’s point that AI agents can be marshaled in a coordinated, adaptive manner, operating as a unified force across platforms and contexts.
Several studies explicitly demonstrate the coordination and influence maximization capabilities of AI swarms. De Marzo et al. [37] find that AI agents can coordinate collective actions at scales beyond human limits, highlighting the potential for swarms to carry out synchronized campaigns (e.g., flooding social networks) with super-human efficiency. Likewise, Li et al. [38] propose an evolutionary agent-based approach to influence maximization in dynamic social networks, illustrating how a swarm of adaptive agents can optimally spread information (or disinformation) through a network by continually learning and targeting influential nodes.
A related framework by Nwala et al. [39] introduces a language-based model for simulating social media account behaviors, which provides a toolkit to craft more realistic and effective bot personas. The effectiveness of such personas is evidenced by Yang et al. [40], who show that fake social media profiles with AI-generated (deepfake) faces are not only common but often perceived as more trustworthy than real profiles. This finding bolsters the observation that AI-generated avatars can greatly enhance a swarm’s ability to blend in and build credibility, thereby amplifying their influence.
Furthermore, an analysis by Yang and Menczer [41] dissects the ’anatomy of an AI-powered malicious social botnet’, confirming that these botnets can combine generative text and imagery to create a web of interlinked, credible-looking agents. Such botnets, once deployed, operate around the clock and can adapt their tactics in vivo, a point underscored by Muennighoff et al. [42] who describe a test-time scaling method that allows agents to rapidly improve their performance during operations. Together, these works provide concrete examples and mechanisms that underpin the main paper’s arguments: AI swarms possess increasingly autonomous, coordinated, and persuasive capabilities far beyond prior manual “bot” networks. This extended literature emphasizes that through advanced training techniques and frameworks, malicious swarms can effectively self-organize, continuously learn (even from minimal feedback), and exploit human social networks with a sophistication and scale that traditional influence operations never had.
S4 Pathways of Harms to Democracy
The pathways by which AI swarms may damage democratic society are further illuminated by a range of empirical studies and real-world observations. In addition to the theoretical pathways outlined in the main text, recent case studies show these harms manifesting in practice. For example, Hung et al. [43] document how ’AI-generated disinformation’ was deployed during Taiwan’s 2024 presidential election and analyze Taiwan’s responses. Their evidence illustrates the “institutional legitimacy erosion” pathway: malicious AI-driven campaigns sought to undermine public trust in electoral outcomes, although timely counter-measures (fact-checking alliances, etc.) helped blunt the impact. This case reinforces the argument that even well-established democracies are now contending with AI-augmented influence attacks. Likewise, experimental work by Goldstein et al. [44] probes how persuasive AI-generated propaganda can be on human subjects. They find that, under certain conditions, AI-crafted propaganda pieces can measurably shift people’s opinions — lending empirical weight to the main article’s warning about “synthetic consensus cascades.” When a swarm amplifies a narrative, the illusion of widespread agreement (even if artificially generated) can sway individuals via social proof, as the propagation dynamics in these studies suggest.
Swarms could also strategically amplify elite incivility or simulate audience enthusiasm for it, creating false social feedback signals that induce politicians to produce more divisive content, compounding polarization [20]. Well before the latest LLM advances, investigations documented how networks of fake accounts could “manufacture reality” on social media by fabricating grassroots support [45]. Swarm-enabled automation is poised to accelerate and cheapen this playbook, deepening platform degradation (“enshittification”) as synthetic consensus crowds out authentic speech. Evidence of ‘LLM Grooming’ in practice. Monitoring groups report that influence operators are already experimenting with “LLM Grooming”—seeding large volumes of duplicative, crawler-targeted content to contaminate future training corpora [29, 47]. These efforts appear optimized for machine consumption (e.g., low human traffic, heavy duplication across domains), consistent with a strategy to calcify fabricated narratives in subsequent model updates.
At the same time, some findings underline the complexity of these harms. Bail et al. [22] assessed the Russian IRA’s influence operation on American Twitter users and found surprisingly limited changes in political attitudes or behaviors as a direct result. This somewhat reassuring result, however, comes with a caveat emphasized in our context: the IRA’s 2017 bots were relatively primitive. The extended capability of AI swarms (e.g., more persuasive content and better targeting) could overcome the resilience observed by Bail and colleagues. In other words, while prior human-curated bot campaigns had only modest effects, the sophisticated, tailored messaging of LLM swarms might achieve what earlier trolls could not — a concern also voiced by Guess et al. [48] who examined the effects of algorithmic versus chronological news feeds on polarization. Guess et al. found that simply altering social media algorithms did not significantly change users’ political attitudes or polarization levels, implying that human biases and social network structures are hard to perturb. However, this in turn suggests that if AI swarms learn to deftly exploit those very biases and network structures (for instance, by micro-targeting individuals with emotionally resonant, false content), the impact could be substantially greater than what benign feed changes or crude bots have achieved. In line with this, Arcila et al. [49] demonstrate a troubling pathway from online hate speech to offline hate crimes. Their study shows that inflammatory online narratives (in this case targeting migrants and minority groups) reliably forecast real-world violence against those groups. This exemplifies the “engineered norm shifts and mass mobilization” pathway discussed in the main paper: AI swarms, by systematically spreading hate or extremist ideologies, can normalize such views and potentially incite some people to violent action.
Another pathway of harm involves subtler shifts in perceptions and discourse. Williams-Ceci et al. [50] find that biased AI writing assistants can gradually shift users’ attitudes on societal issues. This finding supports the article’s concern about “undermined collective intelligence”: if AI systems that many users rely on (for help drafting messages, seeking answers, etc.) inject imperceptible biases or one-sided framings, public opinion can be nudged over time without people realizing the manipulation. Such AI-mediated biasing of individual cognition can aggregate into large-scale opinion shifts, especially when orchestrated by swarms that ensure the biased narratives saturate the infosphere. The conceptual underpinnings of these threats are also being mapped out. Jungherr [51] offers a broad framework for how AI impacts core pillars of democracy. He argues that AI affects self-rule, political equality, information ecosystems, and competition between regimes in distinct ways, many of which align with the harm pathways identified in our paper (e.g., information disorders affecting informed self-rule). Additionally, Summerfield et al. [52] ask “How will advanced AI systems impact democracy?”, and conclude that without intervention, we risk widespread epistemic fragmentation and erosion of democratic norms. Their forward-looking analysis bolsters our argument that emerging AI, especially in the hands of malicious actors, poses systemic risks to public trust and democratic stability.
Finally, classical and recent research on social influence helps explain why AI swarm-driven harms can be so potent. Christakis and Fowler’s book ’Connected’ [53] famously illustrated how opinions and behaviors can spread contagiously through social networks. Airoldi and Christakis [54] provide experimental evidence of social contagion effects in isolated communities, showing that even minor external “nudges” can cascade into significant shifts in collective beliefs. These works underscore a crucial point: democracies rely on the independent formation of opinions, yet both offline social dynamics and orchestrated online swarms can create echo chambers and cascades that undermine that independence. The AI swarms described in the main paper exploit these very dynamics—by injecting coordinated stimuli at scale, they trigger self-reinforcing belief cycles among citizens. Collective accuracy hinges on how many genuinely independent informational signals a group can access before those signals are aggregated. When messages travel through a social network, they often reach multiple members via overlapping paths; this latent redundancy generates informational dependence that inflates confidence without adding new evidence (for instance, illustrated with “pass-the-parcel” simulations showing belief overshoot under high redundancy) [55]. Independence, therefore, functions as an epistemic corrective lens, preserving the diversity of perspectives that the wisdom-of-crowds effect requires while preventing the double-counting that drives shifts to extremity, biased assimilation, and other forms of polarization [56].
S5 Governance Measures and Technical Defenses
Building on the main paper’s discussion of potential safeguards, a rich body of work offers strategies to defend democracy against malicious AI swarms. These span industry initiatives, policy proposals, and technical innovations. For example, Microsoft’s president, Brad Smith [57], in “Meeting the moment: combating AI deepfakes in elections through today’s new tech accord,” calls for a multi-stakeholder agreement to detect and deter deepfakes in political contexts. His advocacy for a tech accord reflects a broader governance strategy: voluntary cooperation between tech firms, governments, and civil society to set standards and share tools against AI-generated deception. On the policy research side, Diakopoulos and Johnson [58] emphasize anticipating ethical implications of AI before they manifest in elections. They argue for proactive measures such as transparency requirements for AI-generated content and clearer legal frameworks to penalize the malicious use of deepfakes. This complements our main paper’s call for forward-looking governance: setting rules and norms in advance to inoculate democracies against swarm-driven disruptions.
A number of surveys diagnose the vulnerabilities in current AI systems, effectively mapping the terrain that governance and technical defenses must cover. Das et al. [59] survey ’security and privacy challenges of large language models,’ and Yao et al. [60] similarly review LLM security. Together, these surveys catalog issues like data poisoning, prompt injection attacks, privacy leaks, and model misuse. By highlighting how LLMs can be attacked or co-opted, they inform the development of both platform-level defenses (e.g., robust model training and validation to reduce vulnerabilities) and policy responses (e.g, guidelines for red-teaming models before deployment). In essence, they support the main article’s implication that technical hardening of AI systems is a necessary component of defense: more secure models would be less prone to malicious manipulation (such as being turned into spam/spread engines or having toxic outputs extracted). On the flip side, other works propose using AI for defense. Shirado and Christakis (2020) demonstrated that introducing benevolent autonomous agents into online groups can increase cooperation and trust [61]. This suggests a possible intervention where “good bots” are deployed to counteract malicious ones — for instance, bots that communicate factual information or mediate to defuse conflicts could help inoculate communities against swarm-fueled falsehoods. Along these lines, Lai et al. (2024) explore how AI collectives might be harnessed to enhance human diversity of thought and enable self-regulation in online discourse [62]. Such approaches flip the script on swarms: rather than solely viewing multi-agent AI as a threat, they consider networked AI agents that empower users, uphold truthful information, or filter out manipulative content as part of the solution set.
Establishing norms and standards is another crucial defense strategy. Baronchelli [63] advocates for “shaping new norms for AI” at the international level, implying that just as we developed norms against cyberattacks or chemical weapons, we need consensus-driven norms (and perhaps treaties) against information warfare by AI. Hall [64] echoes this in her “Governing AI for humanity” report to the United Nations. She emphasizes global coordination, transparency, and human-centric design in AI governance — principles that, if operationalized, would make it harder for malignant swarms to thrive. For instance, transparency mandates (like labeling AI-generated content) and digital provenance standards could reduce a swarm’s ability to pass fabricated content as authentic grassroots opinion. Technical measures are equally important: Chen et al. [65] show that learning from natural language feedback (a form of reinforcement learning from human feedback) can make AI systems align better with human values. Broadly applied, such alignment techniques can curtail the tendency of language models to produce extremist or false outputs, thereby neutering one tool that malicious actors would otherwise exploit. Menczer et al. [66] take a more targeted view in “Addressing the harms of AI-generated inauthentic content.”. They call for a combination of interventions: improving detection algorithms for AI-generated media, disrupting the economics of online disinformation (e.g., through ad network policies), and educating users to recognize automated propaganda. This aligns with our article’s platform-centric defenses (such as better bot detection and content authentication) and societal resilience (like media literacy campaigns). Where content judgments are unavoidable, provenance-first enforcement can be complemented by empirically validated tools: professional fact-checkers show high cross-outlet agreement and accuracy [67]; identity-based, non-confrontational corrections reduce partisan misinformation sharing [68]; and crowdsourced annotations such as Community Notes increase trust in fact-checking on social platforms [69]. Platforms can also weight ranking by source quality signals. Evidence shows that simple, crowdsourced judgments of news source quality reliably track expert assessments and can be used to elevate trustworthy outlets and reduce misinformation exposure [70].
Complementary to human pre-bunking, “model immunization” aims to build resistance directly into AI systems. By proactively training or fine-tuning models to recognize and resist generating or amplifying specific classes of harmful or false content (e.g., via adversarially curated counterexamples, safety-tuned objectives, and evaluation suites), platforms can harden the underlying models that shape information flows [71]. This approach operationalizes inoculation at the model layer, reducing downstream propagation even when malicious inputs evade upstream filters.
The importance of resilient civic institutions is highlighted by Roberts [72], who examined how citizens adapt to state censorship in authoritarian contexts. Her findings — that determined users route around censorship and seek alternative information channels — suggest that purely heavy-handed approaches (like outright banning accounts or content) might be insufficient or even counter-productive. Instead, democratic defenders might focus on resilience: ensuring that truthful information and corrective contexts spread at least as effectively as lies. In fact, Weyl et al. [73] propose a redesign of social media toward “Prosocial Media” architectures. They advocate for systems that use decentralized identity verification (to distinguish real humans from bots) and novel incentive structures that reward constructive discourse. By drastically reducing anonymity and amplifying good-faith actors, such measures could directly counter the creation of “faux publics” by AI swarms (a threat highlighted in our main text).
Supplementary References
  • [1] European Parliament and Council of the European Union, Regulation (EU) 2024/1689 laying down harmonised rules on artificial intelligence (Artificial Intelligence Act) (2024), http://data.europa.eu/eli/reg/2024/1689/oj, official Journal of the European Union, L, 2024/1689.
  • [2] L. Hammond et al., Multi-Agent Risks from Advanced AI (2025), arXiv preprint, arXiv:2502.14143. doi:10.48550/arXiv.2502.14143.
  • [3] E. Pariser, The Filter Bubble: What the Internet Is Hiding from You. Penguin Press (2011).
  • [4] C. E. Robertson, K. S. Del Rosario, J. J. Van Bavel, “Inside the funhouse mirror factory: How social media distorts perceptions of norms,” Current Opinion in Psychology 60, 101918 (2024), doi:10.1016/j.copsyc.2024.101918.
  • [5] S. Vosoughi, D. Roy, S. Aral, “The spread of true and false news online,” Science 359(6380), 1146–1151 (2018), doi:10.1126/science.aap9559.
  • [6] S. Milli, M. Carroll, Y. Wang, S. Pandey, S. Zhao, A. D. Dragan, “Engagement, user satisfaction, and the amplification of divisive content on social media,” PNAS Nexus 4(3), pgaf062 (2025), doi:10.1093/pnasnexus/pgaf062.
  • [7] J. Y. Lee, “The effects of social media criticism against public health institutions on trust, emotions, and social media engagement,” Proceedings of the National Academy of Sciences 122(26), e2422890122 (2025), doi:/10.1073/pnas.2422890122.
  • [8] A. Diel, T. Lalgi, I. C. Schröter, K. F. MacDorman, M. Teufel, A. Bäuerle, “Human performance in detecting deepfakes: A systematic review and meta-analysis of 56 papers,” Computers in Human Behavior Reports 16, 100538 (2024), doi:10.1016/j.chbr.2024.100538.
  • [9] D. M. J. Lazer, M. A. Baum, Y. Benkler, A. J. Berinsky, K. M. Greenhill, F. Menczer, M. J. Metzger, B. Nyhan, G. Pennycook, D. Rothschild, M. Schudson, S. A. Sloman, C. R. Sunstein, E. A. Thorson, D. J. Watts, J. L. Zittrain, “The science of fake news,” Science 359(6380), 1094–1096 (2018), doi:10.1126/science.aao2998.
  • [10] B. Anderson, Imagined Communities: Reflections on the Origin and Spread of Nationalism, in The New Social Theory Reader, pp. 282–288 (Routledge, 2020).
  • [11] J. Habermas, The Structural Transformation of the Public Sphere: An Inquiry into a Category of Bourgeois Society, T. Burger, transl. (MIT Press, Cambridge, MA, 1989).
  • [12] J. B. Bak-Coleman, M. Alfano, W. Barfuss, C. T. Bergstrom, M. A. Centeno, I. D. Couzin, J. F. Donges, M. Galesic, A. Gersick, J. Jacquet, et al., “Stewardship of global collective behavior,” Proceedings of the National Academy of Sciences 118(27), e2025764118 (2021), doi:10.1073/pnas.2025764118.
  • [13] S. Bradshaw, P. N. Howard, The Global Disinformation Order: 2019 Global Inventory of Organised Social Media Manipulation, CyberTroop Report No. 19 (Oxford Internet Institute, 2019), available at https://demtech.oii.ox.ac.uk/wp-content/uploads/sites/93/2019/09/CyberTroop-Report19.pdf [Accessed 22-09-2025].
  • [14] D. Arnaudo, Computational Propaganda in Brazil: Social Bots during Elections, Working Paper No. 2017.8 (Computational Propaganda Research Project, Oxford Internet Institute, 2017), pp. 1–39.
  • [15] #ShePersisted, Big Tech and the Weaponization of Misogyny in the Brazil Online Ecosystem, Tech. Rep. (ShePersisted, 2024), ShePersisted-Brazil Report, available at https://she-persisted.org/wp-content/uploads/2024/04/ShePersisted-Brazil-Report-ENG.pdf [Accessed 22-09-2025].
  • [16] J. C. Ong, J. V. Cabañes, Architects of Networked Disinformation: Behind the Scenes of Troll Accounts and Fake News Production in the Philippines (University of Massachusetts Amherst, 2018), doi:10.7275/2cq4-5396.
  • [17] Rappler Research Team, with Internews, Patient Zero: A Study on the Philippine Information Ecosystem (Rappler/Internews, Feb. 2022), available at: https://internews.org/wp-content/uploads/2022/02/Rappler-Internews-Patient-Zero-A-study-on-the-Philippine-information-ecosystem.pdf [Accessed 22-09-2025].
  • [18] J. Yu, E. Domahidi, D. Gamannossi degl’Innocenti, F. Zollo, “Crisis, Country, and Party Lines: Politicians’ Misinformation Behavior and Public Engagement” (2025), arXiv preprint, arXiv:2502.15321.
  • [19] J. Lasser, S. T. Aroyehun, F. Carrella, A. Simchon, D. Garcia, S. Lewandowsky, “From Alternative Conceptions of Honesty to Alternative Facts in Communications by US Politicians,” Nature Human Behaviour 7, 2140–2151 (2023), doi:10.1038/s41562-023-01691-w.
  • [20] J. A. Frimer, H. Aujla, M. Feinberg, L. J. Skitka, K. Aquino, J. C. Eichstaedt, R. Willer, “Incivility Is Rising Among American Politicians on Twitter,” Social Psychological and Personality Science 14, 259–269 (2023), doi:10.1177/19485506221083811.
  • [21] H. Allcott, M. Gentzkow, “Social Media and Fake News in the 2016 Election,” Journal of Economic Perspectives 31(2), 211–236 (2017), doi:10.1257/jep.31.2.211.
  • [22] C. A. Bail, B. Guay, E. Maloney, A. Combs, D. S. Hillygus, F. Merhout, D. Freelon, A. Volfovsky, “Assessing the Russian Internet Research Agency’s Impact on the Political Attitudes and Behaviors of American Twitter Users in Late 2017,” Proceedings of the National Academy of Sciences 117(1), 243–250 (2020), doi:10.1073/pnas.1906420116.
  • [23] D. J. Ruck, N. M. Rice, J. Borycz, R. A. Bentley, “Internet Research Agency Twitter Activity Predicted 2016 US Election Polls,” First Monday 24(7) (2019), doi:10.5210/fm.v24i7.10107.
  • [24] P. Pomerantsev, This Is Not Propaganda: Adventures in the War Against Reality, 1st ed. (PublicAffairs, 2019), ISBN 978-1541762138.
  • [25] P. Pomerantsev, M. Weiss, The Menace of Unreality: How the Kremlin Weaponizes Information, Culture, and Money (Institute of Modern Russia / The Interpreter, 2014).
  • [26] A. Simchon, W. J. Brady, J. J. Van Bavel, “Troll and Divide: The Language of Online Polarization,” PNAS Nexus 1(1), article pgac019 (2022), doi:10.1093/pnasnexus/pgac019.
  • [27] P. Mozur, “A Genocide Incited on Facebook, With Posts From Myanmar’s Military,” The New York Times (15 Oct. 2018), available at: https://www.nytimes.com/2018/10/15/technology/myanmar-facebook-genocide.html [Accessed 16-05-2025].
  • [28] T. Schipper, “Disinformation by Design: Leveraging Solutions to Combat Misinformation in the Philippines’ 2025 Election,” Data & Policy 7, e39 (2025), doi:10.1017/dap.2025.18.
  • [29] Edelman, 2025 Edelman Trust Barometer (Edelman, 2025), available at: https://www.edelman.com/trust/2025/trust-barometer [Accessed 22-09-2025].
  • [30] M. Brenan, “Americans’ Trust in Media Remains Near Trend Low,” Gallup, Tech. Rep. (2024), available at: https://news.gallup.com/poll/651977/americans-trust-media-remains-trend-low.aspx [Accessed 22-09-2025].
  • [31] B. Kennedy, A. Tyson, “Americans’ Trust in Scientists, Positive Views of Science Continue to Decline,” Pew Research Center (14 Nov. 2023), available at: https://www.pewresearch.org/science/2023/11/14/americans-trust-in-scientists-positive-views-of-science-continue-to-decline/ [Accessed 22-09-2025].
  • [32] Pew Research Center, “Public Trust in Government: 1958–2024,” Pew Research Center, Tech. Rep. (24 June 2024), available at: https://www.pewresearch.org/politics/2024/06/24/public-trust-in-government-1958-2024 [Accessed 22-09-2025].
  • [33] Z. Xi, W. Chen, X. Guo, W. He, Y. Ding, B. Hong, M. Zhang, J. Wang, S. Jin, E. Zhou, R. Zheng, X. Fan, X. Wang, L. Xiong, Y. Zhou, W. Wang, C. Jiang, Y. Zou, X. Liu, Z. Yin, S. Dou, R. Weng, W. Qin, Y. Zheng, X. Qiu, X. Huang, T. Gui, et al., “The Rise and Potential of Large Language Model Based Agents: A Survey,” Science China Information Sciences 68, 121101 (2025), doi:10.1007/s11432-024-4222-0.
  • [34] J. Wei, X. Wang, D. Schuurmans, M. Bosma, B. Ichter, F. Xia, E. H. Chi, Q. V. Le, D. Zhou, “Chain-of-Thought Prompting Elicits Reasoning in Large Language Models,” Advances in Neural Information Processing Systems 35, 24824–24837 (2022), arXiv:2201.11903.
  • [35] F. F. Xu, Y. Song, B. Li, Y. Tang, K. Jain, M. Bao, Z. Z. Wang, X. Zhou, Z. Guo, M. Cao, M. Yang, H. Y. Lu, A. Martin, Z. Su, L. Maben, R. Mehta, W. Chi, L. Jang, Y. Xie, S. Zhou, G. Neubig, “TheAgentCompany: Benchmarking LLM Agents on Consequential Real World Tasks,” arXiv preprint arXiv:2412.14161 (Dec. 2024).
  • [36] S. Hong, M. Zhuge, J. Chen, X. Zheng, Y. Cheng, C. Zhang, J. Wang, Z. Wang, S. K. S. Yau, Z. Lin, L. Zhou, C. Ran, L. Xiao, C. Wu, J. Schmidhuber, “MetaGPT: Meta Programming for a Multi-Agent Collaborative Framework,” arXiv preprint arXiv:2308.00352 (2024).
  • [37] G. De Marzo, C. Castellano, D. Garcia, “Large Language Model Agents Can Coordinate Beyond Human Scale,” arXiv preprint arXiv:2409.02822 (2025).
  • [38] W. Li, Y. Hu, S. Wu, Q. Bai, E. Lai, “ABEM: An Adaptive Agent-Based Evolutionary Approach for Influence Maximization in Dynamic Social Networks,” Applied Soft Computing 136, 110062 (2023), doi:10.1016/j.asoc.2023.110062.
  • [39] A. C. Nwala, A. Flammini, F. Menczer, “A Language Framework for Modeling Social Media Account Behavior,” EPJ Data Science 12, article 33 (2023), doi:10.1140/epjds/s13688-023-00410-9.
  • [40] K.-C. Yang, D. Singh, F. Menczer, “Characteristics and Prevalence of Fake Social Media Profiles with AI-Generated Faces,” Journal of Online Trust and Safety 2(4), 197 (2024), doi:10.54501/jots.v2i4.197.
  • [41] K.-C. Yang, F. Menczer, “Anatomy of an AI-Powered Malicious Social Botnet,” Journal of Quantitative Description: Digital Media 4, 1–36 (2024), doi:10.51685/jqd.2024.icwsm.7.
  • [42] N. Muennighoff, Z. Yang, W. Shi, X. Li, L. Fei-Fei, H. Hajishirzi, L. Zettlemoyer, P. Liang, E. Candès, T. Hashimoto, “s1: Simple Test-Time Scaling,” arXiv preprint arXiv:2501.19393 (2025).
  • [43] C. Hung, W.-C. Fu, C.-C. Liu, H.-J. Tsai, “AI Disinformation Attacks and Taiwan’s Responses During the 2024 Presidential Election,” Graduate Institute of Journalism, National Taiwan University (2024), available at: https://www.thomsonfoundation.org/media/268943/ai_disinformation_attacks_taiwan.pdf [Accessed 22-09-2025].
  • [44] J. A. Goldstein, J. Chao, S. Grossman, A. Stamos, M. Tomz, “How persuasive is AI-generated propaganda?” PNAS Nexus 3, pgae034 (2024), doi:10.1093/pnasnexus/pgae034.
  • [45] N. Gutierrez, “Fake accounts, manufactured reality on social media,” Rappler (3 Oct. 2016), https://www.rappler.com/newsbreak/investigative/148347-fake-accounts-manufactured-reality-social-media/ [Accessed 22-09-2025].
  • [46] American Sunlight Project and EU DisinfoLab, “LLM Grooming: A New Strategy to Weaponise AI for FIMI Purposes,” Webinar by EU DisinfoLab and The American Sunlight Project (10 Apr. 2024), https://www.disinfo.eu/outreach/our-webinars/10-april-llm-grooming-a-new-strategy-to-weaponise-ai-for-fimi-purposes/ [Accessed 22-09-2025].
  • [47] American Sunlight Project, “New Report: Russian Propaganda May Be Flooding AI Models” (23 May 2024), https://www.americansunlight.org/updates/new-report-russian-propaganda-may-be-flooding-ai-models [Accessed 22-09-2025].
  • [48] A. M. Guess, N. Malhotra, J. Pan, P. Barberá, H. Allcott, et al., “How Do Social Media Feed Algorithms Affect Attitudes and Behavior in an Election Campaign,” Science 381(6656), 398–404 (2023), doi:10.1126/science.abp9364.
  • [49] C. Arcila Calderón, P. Sánchez Holgado, J. Gómez, M. Barbosa, H. Qi, A. Matilla, P. Amado, A. Guzmán, D. López-Matías, T. Fernández-Villazala, et al., “From Online Hate Speech to Offline Hate Crime: The Role of Inflammatory Language in Forecasting Violence Against Migrant and LGBT Communities,” Humanities and Social Sciences Communications 11(1), 1–14 (2024), doi:10.1057/s41599-024-03899-1.
  • [50] S. Williams–Ceci, A. Bhat, K. Kadoma, L. Zalmanson, M. Naaman, et al., “Biased AI Writing Assistants Shift Users’ Attitudes on Societal Issues,” PsyArXiv Preprint (2025), doi:10.31234/osf.io/mhjn6_v3.
  • [51] A. Jungherr, “Artificial Intelligence and Democracy: A Conceptual Framework,” Social Media + Society 9(3), article 20563051231186353 (2023), doi:10.1177/20563051231186353.
  • [52] C. Summerfield, L. Argyle, M. Bakker, T. Collins, E. Durmus, T. Eloundou, I. Gabriel, D. Ganguli, K. Hackenburg, G. Hadfield, et al., “How Will Advanced AI Systems Impact Democracy?,” arXiv preprint arXiv:2409.06729 (2024), doi:10.48550/arXiv.2409.06729.
  • [53] N. A. Christakis and J. H. Fowler, Connected: The Surprising Power of Our Social Networks and How They Shape Our Lives. Little, Brown Spark (2009), ISBN 978–0316036146.
  • [54] E. M. Airoldi and N. A. Christakis, “Induction of Social Contagion for Diverse Outcomes in Structured Experiments in Isolated Villages,” Science 384(6695), article eadi5147 (2024), doi:10.1126/science.adi5147.
  • [55] U. Hahn, “Individuals, Collectives, and Individuals in Collectives: The Ineliminable Role of Dependence,” Perspectives on Psychological Science 19(3) (2024), doi:10.1177/17456916231198479.
  • [56] T. D. Pilditch, U. Hahn, and D. Lagnado, “The Problem of Dependency,” Synthese 205(4), article 143 (2025), doi:10.1007/s11229-025-04969-w.
  • [57] B. Smith, “Meeting the Moment: Combating AI Deepfakes in Elections Through Today’s New Tech Accord,” Microsoft On the Issues (16 Feb. 2024), https://blogs.microsoft.com/on-the-issues/2024/02/16/ai-deepfakes-elections-munich-tech-accord/ [Accessed 22-09-2025].
  • [58] N. Diakopoulos and D. Johnson, “Anticipating and Addressing the Ethical Implications of Deepfakes in the Context of Elections,” New Media & Society 23(7), 2072–2098 (2021), doi:10.1177/1461444820925811.
  • [59] B. C. Das, M. H. Amini, and Y. Wu, “Security and Privacy Challenges of Large Language Models: A Survey,” ACM Computing Surveys 57(6), article 3712001 (2025), doi:10.1145/3712001.
  • [60] Y. Yao, J. Duan, K. Xu, Y. Cai, Z. Sun, and Y. Zhang, “A Survey on Large Language Model (LLM) Security and Privacy: The Good, the Bad, and the Ugly,” High-Confidence Computing 4(2), article 100211 (2024), doi:10.1016/j.hcc.2024.100211.
  • [61] H. Shirado and N. A. Christakis, “Network Engineering Using Autonomous Agents Increases Cooperation in Human Groups,” iScience 23(9), article 101438 (2020), doi:10.1016/j.isci.2020.101438.
  • [62] S. Lai, Y. Potter, J. Kim, R. Zhuang, D. Song, and J. Evans, “Evolving AI Collectives to Enhance Human Diversity and Enable Self-Regulation,” arXiv preprint arXiv:2402.12590 (2024), doi:10.48550/arXiv.2402.12590.
  • [63] A. Baronchelli, “Shaping New Norms for AI,” Philosophical Transactions of the Royal Society B: Biological Sciences 379(1897), article 20230028 (2024), doi:10.1098/rstb.2023.0028.
  • [64] D. Hall, Governing AI for Humanity: Final Report of the UN High-Level Advisory Body on Artificial Intelligence. United Nations (Sept. 2024), https://www.un.org/sites/un2.un.org/files/governing_ai_for_humanity_final_report_en.pdf [Accessed 22-09-2025].
  • [65] A. Chen, J. Scheurer, J. A. Campos, T. Korbak, J. S. Chan, S. R. Bowman, K. Cho, and E. Perez, “Learning from Natural Language Feedback,” Transactions on Machine Learning Research (Feb. 2024), https://openreview.net/forum?id=xo3hI5MwvU.
  • [66] F. Menczer, D. Crandall, Y.-Y. Ahn, and A. Kapadia, “Addressing the Harms of AI-Generated Inauthentic Content,” Nature Machine Intelligence 5(7), 679–680 (2023), doi:10.1038/s42256-023-00690-w.
  • [67] S. Lee, A. Xiong, H. Seo, and D. Lee, “‘Fact-checking’ Fact Checkers: A Data-Driven Approach,” Harvard Kennedy School Misinformation Review 4(5) (Oct. 2023), doi:10.37016/mr-2020-126.
  • [68] C. Pretus, A. M. Javeed, D. Hughes, K. Hackenburg, M. Tsakiris, O. Vilarroya, and J. J. Van Bavel, “The Misleading Count: An Identity-Based Intervention to Counter Partisan Misinformation Sharing,” Philosophical Transactions of the Royal Society B: Biological Sciences 379(1897), article 20230040 (2024), doi:10.1098/rstb.2023.0040.
  • [69] C. P. Drolsbach, K. Solovev, and N. Pröllochs, “Community Notes Increase Trust in Fact-Checking on Social Media,” PNAS Nexus 3(7), article pgae217 (2024), doi:10.1093/pnasnexus/pgae217.
  • [70] G. Pennycook and D. G. Rand, “Fighting Misinformation on Social Media Using Crowdsourced Judgments of News Source Quality,” Proceedings of the National Academy of Sciences 116(7), 2521–2526 (2019), doi:10.1073/pnas.1806781116.
  • [71] S. Raza, R. Qureshi, M. Lotif, A. Chadha, D. Pandya, and C. Emmanouilidis, “Just as Humans Need Vaccines, So Do Models: Model Immunization to Combat Falsehoods,” arXiv preprint arXiv:2505.17870 (2025), doi:10.48550/arXiv.2505.17870.
  • [72] M. E. Roberts, “Resilience to Online Censorship,” Annual Review of Political Science 23(1), 401–419 (2020), doi:10.1146/annurev-polisci-050718-032837.
  • [73] E. G. Weyl, L. Thorburn, E. de Keulenaar, J. Mchangama, D. Siddarth, and A. Tang, “Prosocial Media,” arXiv preprint arXiv:2502.10834 (2025).